# Atomic clock interferometry using optical tweezers

### v1

@@ III Tweezer clock interferometry
The amplitude of the expected oscillating interference is determined by $\epsilon$, which gives us a straightforward way to extract this observable. Importantly, for small redshifts ($T\epsilon\ll 1$), the visibility scales linearly with $\epsilon$: $\mathcal{V}\approx T\epsilon/2$. This should be contrasted with a similar interferometric sequence, only without the $\pi$ pulse in the lower path. In that case, the visibility of the interference pattern scales as $\mathcal{V}\approx 1-(T\epsilon)^{2}/8$. The linear scaling of the visibility in our scheme is crucial to make an experimental observation feasible, as we illustrate below.
@@ IV Feasibility analysis
We now turn to estimate the measurable effect in a realistic experimental scenario. The redshift is
$\epsilon=\frac{\omega_{0}}{c^{2}}\Delta\varphi=\frac{\omega_{0}}{c^{2}}mg\Delta h\,\,,$ (17)
@@
with $\Delta h$ being the height difference between the two interferometer arms. Since the effect is linear in atomic mass, it is beneficial to use a relatively massive species. Therefore, we take the atom to be ${}^{171}$b, with the clock transition being ${}^{1}S_{0}-{}^{3}P_{0}$. The energy difference between the clock states corresponds to optical emission with a wavelength of approximately $\lambda=578nm$. The magic wavelength of the tweezer trap for this atom is around $\lambda_{\text{magic}}=759nm$ [46].
Similar to Ref. [2], we assume a separation between the two tweezer arms of 10 mm, aligned in the same direction as Earth’s gravitational acceleration, $g$. Taking a phase accumulation duration of $T=10s$, we obtain a visibility of $\mathcal{V}\approx 0.02$. The oscillating signal appears on top of a background signal of 0.5. Assuming a $1/\sqrt{N}$ scaling of the noise with the number of measurements, it follows that more than $N>625$ measurements are needed to achieve a signal-to-noise ratio better than one. Factoring in an overhead of 5 seconds for each run, the total duration per run extends to 15 seconds. In the worst case scenario where only a single atom is utilized per run, it would take approximately 2.5 hours of data collection for the gravitational time dilation effect to become discernible. Given that the interferometer allows for multiple atoms per run, the effective runtime to achieve a specified signal-to-noise ratio is expected to be significantly reduced.
@@
Operating two optical Ramsey-like pulses coherently over a span of 10 seconds implies a requirement for a narrow linewidth of the clock laser. This requirement is similar to what is needed in light-pulse interferometers using the optical transition in ${}^{88}$r [47]. Indeed, state-of-the-art laser systems built for atomic clocks can reach a linewidth below 10 mHz [48], achieving coherence times that exceed 10 seconds [29]. For the estimation of the redshift measurement previously described, such a laser is sufficient.
An alternative approach is to use a commercially available laser with a larger linewidth, but prolong its coherence time using the Yb atoms. This can be done, for example, through phase-locking the laser to the atomic clock transition with quantum non-demolition (QND) measurements [49]. Such methods were able to achieve clock coherence time on the order of 1-5 seconds. Interferometry time of 1 second will result in a visibility $\mathcal{V}\approx 0.002$. Assuming the same overhead time, this will require integrating data over approximately 11 days when working with a single atom, or about a day, if 10 atoms are used in each run. Either way, we conclude that the observation of gravitational redshift with ACIF is within the reach of current technological capabilities.
@@ V Discussion
We introduce a guided atomic clock interferometer approach, an extension of our previous tweezer interferometer method [2], tailored for atoms with two internal states. This technique employs optical laser pulses to create a superposition of the internal states and utilizes precise manipulation of the tweezers’ position and intensity for the spatial adiabatic splitting and recombining of the wave packet. After completion of the interferometric scheme, we record the population within each clock state and exit port. The coherent oscillations observed between the two exit ports confirm the wave packet’s spatial coherence throughout the experiment. Oscillations between the internal states reveal time dilation across the paths. In particular, the visibility of these oscillations is instrumental in determining the gravitational redshift between the interferometer arms. Clock interferometry, which has not been performed to date, is achievable using our proposed scheme with the current state of technology.
Quantum theory has been tested successfully in many experiments, including with atomic interferometry. Similarly, predictions of General Relativity have also been verified, including measurements of the gravitational redshift using two separate clocks [12, 13, 14, 15, 16, 17]. However, theories of quantum gravity in regimes in which both quantum mechanics and general relativity are relevant remain untested. The proposed ACIF could probe this unexplored regime for the first time [18].
@@ Acknowledgements.
We thank Jonathan Nemirovsky and Eliahu Cohen for helpful discussions. We thank Anastasiya Vainbaum and Amir Stern for critical reading of the manuscript. This research was supported by the Israel Science Foundation (ISF), grant No. 3491/21, and by the Pazy Research Foundation. This research project was partially supported by the Helen Diller Quantum Center at the Technion.

### v3

@@ III Tweezer clock interferometry
Since this amplitude is determined by $\epsilon$, it gives a straightforward way to extract its value. Importantly, for small redshifts ($T\epsilon\ll 1$), the visibility scales favorably linearly with $\epsilon$: $\mathcal{V}\approx T\epsilon/2$. This should be contrasted with a similar interferometric sequence, only without the $\pi$ pulse in the lower path. In that case, the visibility of the interference pattern scales as $\mathcal{V}\approx 1-(T\epsilon)^{2}/8$. The linear scaling of the visibility in our scheme is crucial to make an experimental observation feasible, as we illustrate below.
@@ IV Feasibility analysis
We first estimate the measurable effect in a realistic experimental scenario. The redshift is
$\epsilon=\frac{\omega_{0}}{c^{2}}gh\,\,,$ (18)
@@
with $h$ being the height difference between the two interferometer arms. We take the atom to be 171Yb, with the clock transition being ${}^{1}S_{0}-{}^{3}P_{0}$. The energy difference between the clock states corresponds to optical emission with a wavelength of approximately $\lambda=578$nm. The magic wavelength of the tweezer trap for this atom is around $\lambda_{\text{magic}}=759$nm [51].
The resilience of the measurement to differential intensity fluctuations depends on the stability of the tweezer wavelength. Deviations from the magic wavelength result in intensity-dependent light-shifts in the clock states frequencies [52, 53], which in turn may reduce the accuracy. The tweezers’s wavelength can be stabilized to the desired value with sub-MHz resolution [51, 54], resulting in frequency noise in the order of $10^{-5}$z. This noise level has no effect on the measurement accuracy, as we have confirmed through numerical simulations.
To minimize motional transitions in the tweezer during the optical pulse, it is required to operate within the Lamb-Dicke regime [32]. This regime is characterized by a small $\eta=\frac{2\pi}{\lambda}\cdot x_{0}$, where $x_{0}$ represents the spatial extent of the atomic wavefunction trapped in the tweezer. For a tweezer with a depth of $300\mu$K and a waist of $1\mu$m, the resulting Lamb-Dicke parameters are approximately $\eta\approx 0.3$ and $0.73$ in the radial and axial directions, respectively. These values fall within the operational range for optical atomic clocks [34], suggesting they should be adequate for the interferometer. During the phase accumulation stage, in the absence of optical pulses, a deep trap is unnecessary. To lower the chance of spontaneous emission from the tweezer light during this extended phase, the depth of the tweezers should be reduced and then increased again just before the recombination stage.
Similar to Ref. [2], we assume a separation between the two tweezer arms of $h=10$ mm, aligned in the same direction as Earth’s gravitational acceleration, $g$. Taking a phase accumulation duration of $T=10s$, we obtain a visibility of $\mathcal{V}\approx 0.02$. The oscillating signal appears on top of a background signal of 0.5. To estimate the number of runs required to clearly observe the gravitational effect, we have performed a Monte Carlo simulation of the entire experiment. In the simulation, we set the detuning to $\Delta=2\pi\cdot 1000$ Hz and scan $N_{1}$ different durations in the range $T\in[10,10+\frac{2\pi}{\Delta}]$. For each of these durations, we simulate the measurement with $N_{a}$ atoms according to Eqs. 15 and 16. We then repeat the procedure $N_{2}$ times for each duration to get an estimator of $P_{g}(T)$ and $P_{1}(T)$. We fit the fringe of the former, and from the extracted visibility, we find $\epsilon$. Additionally, we included intensity fluctuations between the tweezer arms by adding a random energy shift on the order of $\hbar\Delta$ between the $E_{g;2}$ and $E_{g;1}$ states.
A typical result of the simulation is shown in figure 2. In this example, we have taken $N_{a}=20$ atoms per run and $N_{2}=5000$ repetitions. Factoring in an overhead of 5 seconds for each run, the total duration per run extends to 15 seconds. Scanning the fringe at $N_{1}=8$ different phases, as shown in the figure, requires about a week of data collection. With these parameters, the fringe contrast is clearly visible, and the relative accuracy (one standard deviation) in determining $\epsilon$ is approximately $9\%$. Importantly, the accuracy of the extracted $\epsilon$ is not compromised by the intensity fluctuations between the tweezers. In Table 1, we present the relative accuracy for other realistic choices of parameters. In all cases, the accuracy allows unambiguous determination of a gravitational redshift effect.
Figure 2: Numerical simulation of the interferometric measurement with a separation of $h=10$ mm in Earth’s gravitational field. In this simulation, we used $\Delta=2\pi\cdot 1000$ Hz, $N_{a}=20$ atoms per run, and $N_{2}=5000$ repetitions of each duration. The circular markers represent the average of the estimator of $P_{g}$, with error bars that represent the standard error. The solid line is the fit to Equation 16, with the amplitude and phase being the free parameters.
Number of atoms in a run ($N_{a}$) Number of repetitions per phase and duration ($N_{2}$) Probing duration (T) [s] total runtime Relative accuracy in extracting $\epsilon$
100 5000 1 $\sim$ 2.8 days 38.1%
100 10000 1 $\sim$ 5.6 days 28.4%
20 5000 3 $\sim$ 3.7 days 28.9%
100 5000 3 $\sim$ 3.7 days 12.9%
100 10000 3 $\sim$ 7.4 days 9.8%
10 1000 10 $\sim$ 1.4 days 28.4%
10 5000 10 $\sim$ 7 days 12.4%
20 5000 10 $\sim$ 7 days 8.8%
100 1000 10 $\sim$ 1.4 days 9%
100 5000 10 $\sim$ 7 days 3.8%
100 10000 10 $\sim$ 13.9 days 2.7%
Table 1: Different possible choices of experimental parameters. In all cases, we assume a separation of $h=10$ mm between the interferometer arms in Earth’s gravitational field. Similar to figure 2, we fix the number of different phases to $N_{1}=8$. The total runtime takes into account an overhead of $5$ seconds per run. The relative accuracy is defined as one standard deviation of the extracted values of $\epsilon$ over an ensemble of $1000$ Monte Carlo simulation runs, normalized by the theoretical value of $\epsilon$.
The Monte Carlo simulations allow us to test our approach to verifying the coherent splitting. The random fluctuations we introduce in the relative ground state energies translate into essentially a random phase of the first sine term in Eq. 15. In figure 3, we plot the histogram of the estimator of $P_{1}$ (darker shading) for $10$ (top) and $100$ (bottom) atoms in each run. For comparison, we also depict the histogram obtained if a collapse randomly occurs to one of the two arms (brighter shading). The coherent and incoherent histograms are markedly different. In particular, the probability of finding highly unbalanced splitting between the output ports is negligibly small when the wavepacket decoheres, while it is maximal if it maintains its coherence. These results prove that the coherence of the wavepacket can be verified even in the presence of strong intensity fluctuations and with a relatively small number of atoms.
Figure 3: Histograms of the estimator of $P_{1}$ for coherent and incoherent splitting. The histograms are from an ensemble of 1000 Monte Carlo simulation runs that include strong intensity fluctuations between the tweezer arms. In the incoherent splitting case (brighter shading), we assume that the wavepacket randomly collapses to one of the arms after the splitting stage. For both $N_{a}=10$ atoms (top) and $N_{a}=100$ atoms (bottom) per run, the coherent and incoherent histograms are easily distinguished. The parameters in these simulations are $T=10.00025$ sec, $\Delta=2\pi\cdot 1000$Hz, $h=10$mm, and $N_{2}=5000$. The different discretization of the x-axis is due to the different $N_{a}$, which results in a different finite set of possible values of $P_{1}$ in each run.
@@
Operating two optical Ramsey-like pulses coherently over several seconds requires a narrow linewidth of the clock laser. This requirement is similar to what is needed in light-pulse interferometers using the optical transition in 88Sr [55]. State-of-the-art laser systems built for atomic clocks can achieve a linewidth below 10 mHz [56, 57]. A recent experiment with a tweezer atomic clock reported an atom-light coherence time of 3.6 seconds [34]. The reduced coherence time relative to the limit dictated by the laser linewidth is likely due to fluctuations in the tweezer position, or uncontrolled phase noise from environmental factors. Even so, as shown in Table 1, this time is sufficient to determine the redshift with an accuracy better than $10\%$. We conclude that the observation of gravitational redshift with ACIF is within the reach of current technological capabilities.
@@ V Discussion
We introduce a guided atomic clock interferometer approach, an extension of our previous tweezer interferometer method [2], tailored for atoms with two internal states. This technique employs optical laser pulses to create a superposition of the internal states and utilizes precise manipulation of the tweezers’ position and intensity for the spatial adiabatic splitting and recombining of the wavepacket. After completion of the interferometric scheme, we record the population within each clock state and exit port. The statistical distribution of the splitting between the two exit ports confirms the wavepacket’s spatial coherence throughout the experiment. Oscillations between the internal states reveal time dilation across the paths. In particular, the visibility of these oscillations is instrumental in determining the gravitational redshift between the interferometer arms. Clock interferometry, which has not been performed to date, is achievable using our proposed scheme with the current state of technology.
Quantum theory has been tested successfully in many experiments, including with atomic interferometry. Similarly, predictions of General Relativity have also been verified, including measurements of the gravitational redshift using two separate clocks [16, 17, 18, 19, 20, 21]. However, theories concerning regimes in which both quantum mechanics and general relativity are relevant remain untested. The proposed ACIF could probe this unexplored regime for the first time [22].
@@ Acknowledgements.
We thank Jonathan Nemirovsky, Eliahu Cohen, and Ran Finkelstein for helpful discussions. We thank Anastasiya Vainbaum, Amir Stern, and Magdalena Zych for critical reading of the manuscript. This research was supported by the Israel Science Foundation (ISF), grant No. 3491/21, and by the Pazy Research Foundation. This research project was partially supported by the Helen Diller Quantum Center at the Technion.